# KramaBench: A Benchmark for AI Systems on Data Intensive Tasks

### v2

@@ Abstract
Discovering insights from a real-world data lake potentially containing unclean, semi-structured, and unstructured data requires a variety of data processing tasks, ranging from extraction and cleaning to integration, analysis, and modeling. This process often also demands domain knowledge and project-specific insight. While AI models have shown remarkable results in reasoning and code generation, their abilities to design and execute complex pipelines that solve these data-lake-to-insight challenges remain unclear. We introduce KramaBench 111Assets available at https://github.com/mitdbg/Kramabench which consists of 104 manually curated and solved challenges spanning 1700 files, 24 data sources, and 6 domains. KramaBench focuses on testing the end-to-end capabilities of AI systems to solve challenges which require automated orchestration of different data tasks. KramaBench also features a comprehensive evaluation framework assessing the pipeline design and individual data task implementation abilities of AI systems. Evaluating 8 LLMs with our single-agent reference framework DS-Guru, alongside open- and closed-source agentic systems, we find that while current single-agent systems may handle isolated data-science tasks and generate plausible draft pipelines, they struggle with producing working end-to-end pipelines. On KramaBench, the best system reaches only 50% end-to-end accuracy in the full data-lake setting. Even with perfect retrieval, the accuracy tops out at 59%. Leading LLMs can identify up to 42% of important data tasks but can only fully implement 20% of individual data tasks.
@@ 1 Introduction
While recent research has advanced individual components of these workflows such as code generation (Nam et al., 2024; Wang & Chen, 2023), tool use (Qin et al., 2024b, a), and natural language question answering (Zhang et al., 2024b; Pu et al., 2023), the challenge of designing and executing complete end-to-end data science pipelines remains underexplored.
@@
We evaluated KramaBench across eight models, along with three different configurations of DS-Guru and three other existing agentic systems (Hugging Face, 2025; OpenAI, 2025; Google, 2025). We conducted extensive ablations studies and failure analyses, taking advantage of our comprehensive evaluation framework and obscured inputs.
Through KramaBench, we observed multiple insights about where LLM systems are successful: (1) Agentic control flow is helpful with KramaBench’s challenges: systems like smolagents that iteratively search, plan, and repair achieve 50% end-to-end accuracy, outperforming the strongest configurations of DS-Guru (22% overall), which uses a structured control flow. (2) LLM systems can reason at a coarse level about the data operations required by a successful workflow and generate plausible pipelines, achieving 42% on pipeline design.
Our analyses also reveal some persistent challenges: (1) Retrieval from a data lake is problematic, but not the dominating obstacle. Supplying only the gold files improves overall accuracy by only 9-10% across systems using different retrieval mechanisms. (2) Weaknesses in fine-grained data-dependent reasoning cause models to fail. Systems even fail most of the time at implementing individual simple sub-tasks, capping at 19.75% when evaluated under the individual task implementation described above. (3) Agents often fail to achieve a holistic understanding of the data lake. We observe that the agents often overly rely on their prior knowledge (12%-16% performance fluctuation on obscured inputs), or assume clarifications will be given from a user (22% of failures).
@@ 2 The Design of KramaBench
Tasks in KramaBench are based on real-world data science challenges from six domains: archaeology, astronomy, biomedical research, environmental science, legal insight discovery, and wildfire prevention. Each domain is associated with a data lake containing raw files in structured, semi-structured, or unstructured formats from multiple sources. Each task is a natural language description of a domain-specific data science problem. The goal of a system under test is to design and execute an end-to-end pipeline that takes the entire domain data lake as input and produces the correct output. In addition to the target answer, KramaBench provides the ground truth solution both in code and in annotated sub-tasks: natural language descriptions of smaller building-block operations that are essential elements within a full solution along with a prompt and their target answers. These finer-grained references enable the evaluations of pipeline design and individual task implementation.
@@ 2.1 Task Design and Validation
To curate tasks, we started with published studies and reports that (1) contain quantitative or graphical findings produced by data analysis, (2) are based on complete and publicly accessible datasets, and (3) require complex multi-step pipelines involving heterogeneous and noisy inputs. Grounding onto these studies and reports ensures that our tasks reflect real-world data science pipelines. We followed a 4-step workflow involving tight validations and repeated verifications of reference solutions to ensure the quality of tasks, reference solutions, and fine-grained annotations.
@@
| Archeology | 12 | 71 | 50.00% | 5 | 2 | 7.5MB |
| Astronomy | 12 | 68 | 50.00% | 1556 | 8 | 486MB |
| Biomedical | 9 | 38 | 66.66% | 7 | 2 | 175MB |
| Environment | 20 | 148 | 70.00% | 37 | 3 | 31MB |
| Legal | 30 | 188 | 53.33% | 136 | 2 | 1.3MB |
| Wildfire | 21 | 120 | 71.42% | 23 | 7 | 1GB |
| Total | 104 | 633 | 60.58% | 1764 | 24 | 1.7GB |
Table 2 reports the statistics and difficulty distributions of the 6 domains and their tasks. We provide more detailed descriptions and an example of tasks, key functionalities, and sub-tasks in Appendix D.
@@ 2.2 Evaluation Mechanism
As discussed in Section 1, KramaBench evaluates systems on three capabilities. Our primary focus is (1) end-to-end automation.
(1) End-to-end Automation. For each task, the system output is given a score in based on the reference target answer. The scoring schemes for each possible answer type address fuzzy matches and are discussed in Table 3. Given a domain workload consisting of numerous tasks , the total score of a system for is . The score of for the entire benchmark suite is analogous.
| Type | Example | Metric | Scoring | ||
The name of a file to load. | Accuracy | |||
The month when an experiment started. | ParaPluie paraphrase detection (Lemesle et al., 2025) | |||
Counting the number of entries satisfying a predicate. | Accuracy | |||
Prediction of a future experiment observation. | Relative Absolute Error (RAE) | |||
Names of columns to filter data. | F1 (exact match) | F1 score | ||
Regression coefficients for different variables. | F1 score (approximate match > 0.9) | F1 score |
Results under the following two less-automated evaluation settings provide insights into why a system may succeed or fail in the end-to-end automation setting and the abilities of a system to assist with a human-in-the-loop.
(2) Pipeline Design. For each task, we assess the system generated pipeline using the key functionalities that any correct pipeline needs to contain in some form (from Step 3 of Subsection 2.1). We score the system with the fraction of key functionalities covered in the pipeline produced by the system. Coverage is evaluated via LLM-as-a-judge following the method in Tong & Zhang (2024) using the description obtained in Step 3.
(3) Sub-task Evaluation. We provide systems with the problem statements for sub-tasks and compare system outputs to human-curated target answers (as in Step 4 of Subsection 2.1) using the same scoring approach as in end-to-end evaluation. Full technical details of these evaluations are provided in Appendix E.
@@ 2.3 Reference Implementation
We introduce DS-Guru, a lightweight framework that serves as minimal scaffolding to enable a single out-of-the-box LLM to attempt the data science challenges in KramaBench. DS-Guru has three variants. No-context: The LLM is invoked one-shot with the problem description and the names and paths of the files from the data lake, without any file contents. One-shot: The LLM is invoked one-shot with the problem description and sample snippets from each data file. Few-shot: The LLM is first invoked once with the task description and sample snippets, then re-invoked few-shot with execution results and error messages from the pipeline it implemented in the previous shot. With all variants, DS-Guru instructs the LLM to decompose the task into simpler tasks before attempting to implement each task provide the concrete pipeline implementation along with the answer.
@@ 3 Experimental Setup
We accessed all LLMs in different systems via OpenAI and Together APIs; pipelines generated by systems are executed locally.
DS-Guru: We combine each of the three variants (as in Subsection 2.3) of DS-Guru with six LLMs: GPT-o3, GPT-4o, Claude-3.5-Sonnet, Llama3.3, Deepseek-R1-70B, and Qwen2.5-Coder-32B (OpenAI, 2025, 2024; Anthropic, 2024; Meta AI, 2025; DeepSeek-AI et al., 2025; Hui et al., 2024), totaling to 18 concrete DS-Guru implementations.
smolagents Deep Research (smolagents DR): We evaluate Hugging Face smolagents (Hugging Face, 2025), an open source deep research agentic system, on the single agent setting. We report results with four different LLMs: GPT-o3, GPT-4o, Claude-3.5-Sonnet, and Claude-3.7-Sonnet. We view smolagents DR as a representative system for open-source “deep research” projects (e.g., Alibaba’s Academy (2025)).
Closed-Source Deep Research Systems: OpenAI Deep Research (OpenAI DR, OpenAI (2025)) and Gemini Pro-2.5 Agentic Mode (Gemini Agentic) (Google DeepMind, 2025; Google, 2025) were evaluated manually through their web interfaces under the end-to-end automation setting. We made best efforts instructing them not to search online. However, this restriction was not enforceable.
@@
| smolagents DR | Agentic retrieval | Full, Trimmed, Oracle | Agentic loops | Off |
@@
We evaluated four different systems, which differ in four important ways.
@@
Control flow simply describes whether a system has conventional loops or agentic flow. Internet access describes whether systems have web search capabilities.
We measured the LLM cost of running the full KramaBench evaluation for a system under test. For DS-Guru (few-shot, GPT-o3), a variant with high pipeline design and sub-task implementation completion, evaluating end-to-end answers took 4,501 tokens; evaluating pipeline designs took 116,805 tokens; and evaluating sub-task implementation took 10,358 tokens.
@@ 4 Results and Takeaways
| Archaeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Overall | ||
@@
| DS-Guru no-context | GPT-o3 | 25% | 1.73% | 3.50% | 1.35% | 3.35% | 24.87% | 9.64% |
| GPT-4o | 0.00% | 1.41% | 1.98% | 0.45% | 1.46% | 1.45% | 1.62% | |
| Claude-3.5 | 16.67% | 1.62% | 2.87% | 1.17% | 7.33% | 13.63% | 7.45% | |
| DS-Guru one-shot | GPT-o3 | 25% | 3.00% | 8.63% | 7.66% | 19.15% | 45.95% | 20.80% |
| GPT-4o | 8.33% | 1.40% | 9.38% | 2.60% | 2.74% | 19.39% | 7.61% | |
| Claude-3.5 | 0.00% | 4.15% | 2.15% | 6.21% | 6.68% | 34.99% | 10.85% | |
| DS-Guru few-shot | GPT-o3 | 25% | 3.53% | 8.95% | 19.6% | 13.89% | 50.73% | 22.08% |
| GPT-4o | 16.67% | 2.76% | 8.97% | 2.60% | 2.80% | 17.18% | 8.28% | |
| Claude-3.5 | 16.67% | 1.52% | 1.96% | 11.21% | 7.01% | 39.16% | 14.35% | |
| smolagents DR | GPT-o3 | 41.67% | 16.67% | 33.33% | 50% | 50% | 38.1% | 41.36% |
| GPT-4o | 33.33% | 0.00% | 11.11% | 35% | 40% | 38.1% | 30.77% | |
| Claude-3-5 | 33.33% | 0.00% | 22.22% | 60% | 46.67% | 52.38% | 41.35% | |
| Claude-3-7 | 33.33% | 16.67% | 44.44% | 60% | 63.33% | 52.38% | 50% | |
@@
| DS-Guru no-context | GPT-o3 | 17.83% | 12.93% | 19.48% | 19.17% | 9.94% | 16.13% | 14.93% |
| GPT-4o | 15.09% | 9.15% | 12.16% | 11.26% | 8.88% | 7.15% | 10.05% | |
| Claude-3.5 | 16.52% | 10.63% | 9.87% | 12.51% | 9.80% | 0.00% | 11.63% | |
| DS-Guru one-shot | GPT-o3 | 23.90% | 21.14% | 18.29% | 28.48% | 18.49% | 25.08% | 22.85% |
| GPT-4o | 14.26% | 10.58% | 9.38% | 20.37% | 10.96% | 19.21% | 14.86% | |
| Claude-3.5 | 17.07% | 10.24% | 9.44% | 22.27% | 11.47% | 17.93% | 15.48% | |
| DS-Guru few-shot | GPT-o3 | 27.78% | 23.22% | 19.56% | 33.67% | 35.14% | 32.53% | 31.92% |
| GPT-4o | 18.97% | 19.29% | 12.51% | 27.14% | 25.23% | 26.07% | 23.60% | |
| Claude-3.5 | 16.24% | 14.02% | 14.80% | 33.83% | 26.36% | 25.02% | 24.22% | |
| smolagents DR | GPT-o3 | 41.67% | 25% | 44.44% | 45% | 44.83% | 47.62% | 44.45% |
| GPT-4o | 25% | 25% | 22.22% | 20% | 56.67% | 38.1% | 39% | |
| Claude-3-5 | 16.67% | 25% | 33.33% | 25% | 66.66% | 66.66% | 47% | |
| Claude-3-7 | 41.67% | 33.33% | 77.78% | 80% | 63.33% | 71.43% | 59% | |
@@
| DS-Guru few-shot | GPT-o3 | 25.00% | 3.17% | 2.71% | 17.02% | 16.25% | 49.42% | 21.78% |
| smolagents DR | Claude-3-7 | 33.33% | 33.33% | 44.44% | 65% | 63.33% | 66.67% | 57.85% |
| OpenAI DR | GPT-o3-dr | 40% | 33.33% | 44.45% | 61.67% | 50% | 67.28% | 52.18% |
| Gemini Agentic | Gemini-2.5-Pro | 25% | 16.67% | 33.33% | 25% | 13.33% | 24.87% | 18.48% |
Agentic control flows drive the largest performance gains on KramaBench. smolagents DR (Claude-3-7, max agentic iterations is 20) consistently outperforms DS-Guru across all domains (Table 9), reaching 50% overall score compared to the best DS-Guru variant (few-shot, GPT-o3; 22.08%). The DS-Guru (few-shot), which enables the LLM to catch implementation errors only moderately improves over DS-Guru (one-shot, GPT-o3), with 1.28% overall improvement. Our detailed studies increased few-shot to 20 iterations yet still showed minor improvements (Appendix Table 12). This indicates that despite the heterogeneity of data operations, the core challenges are not isolated data operation implementation issues, but instead are to (1) explore and fix the design choices of the end-to-end pipeline; (2) iteratively understand the data and schema in a large data lake. Smolagent DR’s agentic control flow helps address these challenges. Note that in the Trimmed setting (max 10 files per call), OpenAI DR reaches 52.18% overall, partly due to its web search capability.
In terms of cost, smolagents DR (Claude-3-7) averages 6.10 minutes per task—faster than OpenAI DR (10.35) but more than 10× slower than DS-Guru few-shot (0.76).

### v3

# KramaBench: A Benchmark for AI Systems on Data-to-Insight Pipelines over Data Lakes

@@ Abstract
Discovering insights from a real-world data lake potentially containing unclean, semi-structured, and unstructured data requires a variety of data processing tasks, ranging from extraction and cleaning to integration, analysis, and modeling. This process often also demands domain knowledge and project-specific insight. While AI models have shown remarkable results in reasoning and code generation, their abilities to design and execute complex pipelines that solve these data-lake-to-insight challenges remain unclear. We introduce KramaBench 111Assets available at https://github.com/mitdbg/Kramabench which consists of 104 manually curated and solved challenges spanning 1700 files, 24 data sources, and 6 domains. KramaBench focuses on testing the end-to-end capabilities of AI systems to solve challenges which require automated orchestration of different data tasks. KramaBench also features a comprehensive evaluation framework assessing the pipeline design and individual data task implementation abilities of AI systems. We evaluate 8 LLMs using our single-agent reference framework DS-Guru, alongside both open- and closed-source single- and multi-agent systems, and find that while current agentic systems may handle isolated data-science tasks and generate plausible draft pipelines, they struggle with producing working end-to-end pipelines. On KramaBench, the best system reaches only end-to-end accuracy in the full data-lake setting. Even with perfect retrieval, the accuracy tops out at . Leading LLMs can identify up to of important data tasks but can only fully implement of individual data tasks.
@@ 1 Introduction
While recent research has advanced individual components of these workflows such as code generation (Nam et al., 2024; Wang and Chen, 2023), tool use (Qin et al., 2024b; a), and natural language question answering (Zhang et al., 2024b; Pu et al., 2023), the challenge of designing and executing complete end-to-end data science pipelines remains underexplored.
@@
We evaluated KramaBench across eight models, along with three different configurations of DS-Guru and three other existing agentic systems (Hugging Face, 2025; OpenAI, 2025a; Google, 2025). We conducted extensive ablations studies and failure analyses, taking advantage of our comprehensive evaluation framework and obscured inputs.
Through KramaBench, we observed multiple insights about where LLM systems are successful: (1) Agentic control flow is helpful with KramaBench’s challenges: a canonical single-agent system (smolagents DR) that iteratively search, plan, and repair achieve end-to-end accuracy, outperforming the strongest configurations of DS-Guru ( overall), which uses a structured control flow. (2) a canonical multi-agent system (smolagents Reflexion, Shinn et al. (2023)) using an evaluator agent and reflections achieves end-to-end accuracy. The marginal difference () relative to the single-agent baseline suggests that simple evaluator-style coordination alone does not substantially improve performance on data-intensive workflows, indicating opportunities for more specialized multi-agent designs.(3) LLM systems can reason at a coarse level about the data operations required by a successful workflow and generate plausible pipelines, achieving on pipeline design.
Our analyses also reveal some persistent challenges: (1) Retrieval from a data lake is problematic, but not the dominating obstacle. Supplying only the gold files improves overall accuracy by only across systems using different retrieval mechanisms. (2) Weaknesses in fine-grained data-dependent reasoning cause models to fail. Systems even fail most of the time at implementing individual simple sub-tasks, capping at when evaluated under the individual task implementation described above. (3) Agents often fail to achieve a holistic understanding of the data lake. We observe that the agents often overly rely on their prior knowledge (at most performance fluctuation on obscured inputs), or assume clarifications will be given from a user ( of failures).
@@ 2 The Design of KramaBench
Tasks in KramaBench are based on real-world data science challenges from six domains: archeology, astronomy, biomedical research, environmental science, legal insight discovery, and wildfire prevention. Each domain is associated with a data lake containing raw files in structured, semi-structured, or unstructured formats from multiple sources. Each task is a natural language description of a domain-specific data science problem. The goal of a system under test is to design and execute an end-to-end pipeline that takes the entire domain data lake as input and produces the correct output. In addition to the target answer, KramaBench provides the ground truth solution both in code and in annotated sub-tasks: natural language descriptions of smaller building-block operations that are essential elements within a full solution along with a prompt and their target answers. These finer-grained references enable the evaluations of pipeline design and individual task implementation.
@@ 2.1 Task Design and Validation
To curate tasks, we started with published studies and reports that (1) contain quantitative or graphical findings produced by data analysis, (2) are based on complete and publicly accessible datasets, and (3) require complex multi-step pipelines involving heterogeneous and noisy inputs. Grounding onto these studies and reports ensures that our tasks reflect real-world data science pipelines. We followed a 4-step workflow involving tight validations and repeated verifications of reference solutions to ensure the quality of tasks, reference solutions, and fine-grained annotations. We summarize the process in Figure 2.
@@
| Domain | # Tasks (sub-) | %Hard Tasks | # Files (size) |
| Archeology | 12 (71) | 50.00% | 5 (7.5MB) |
| Astronomy | 12 (68) | 50.00% | 1556 (486MB) |
| Biomedical | 9 (38) | 66.66% | 7 (175MB) |
| Environment | 20 (148) | 70.00% | 37 (31MB) |
| Legal | 30 (188) | 53.33% | 136 (1.3MB) |
| Wildfire | 21 (120) | 71.42% | 23 (1GB) |
| Total | 104 (633) | 60.58% | 1764 (1.7GB) |
| Type | Metric score |
| String (exact) | Accuracy () |
| String (approximate) | ParaPluie score () |
| Numeric (exact) | Accuracy () |
| Numeric (approximate) | (0-1) |
| List (exact) | F1 score (0-1) |
| List (approximate) | F1 score (if match > 0.9) |
Table 3 reports the statistics and difficulty distributions of the 6 domains and their tasks. We provide more detailed descriptions and an example of tasks, key functionalities, and sub-tasks in Appendix E.
@@ 2.2 Evaluation Mechanism
As discussed in Section 1, KramaBench evaluates systems on three capabilities. In Figure 3, we provide an overview of these 3 areas. Our primary focus is (1) end-to-end automation.
(1) End-to-end Automation. For each task, the system output is given a score in based on the reference target answer. The scoring schemes for each possible answer type address fuzzy matches and are discussed in Table 3. The string approximation metric uses the method introduced in (Lemesle et al., 2025). Our validation study (details in Subsection G.2) for this LLM-as-a-judge method shows 84% agreement between human annotators and the LLM. Given a domain workload consisting of numerous tasks , the total score of a system for is . The score of for the entire benchmark suite is analogous.
Results under the following two less-automated evaluation settings provide insights into why a system may succeed or fail in the end-to-end automation setting and the abilities of a system to assist with a human-in-the-loop. Figure 8 describes the mapping of the following tasks.
(2) Pipeline Design. For each task, we assess the system generated pipeline using the key functionalities that any correct pipeline needs to contain in some form (from Step 3 of Subsection 2.1). We score the system with the fraction of key functionalities covered in the pipeline produced by the system. Coverage is evaluated via LLM-as-a-judge following the method in Tong and Zhang (2024) using the description obtained in Step 3.
(3) Sub-task Evaluation. We provide systems with the problem statements for sub-tasks and compare system outputs to human-curated target answers (as in Step 4 of Subsection 2.1) using the same scoring approach as in end-to-end evaluation. Full technical details of these evaluations are provided in Appendix G.
@@ 2.3 Reference Implementation
We introduce DS-Guru (Figure 4), a lightweight framework that serves as minimal scaffolding to enable a single out-of-the-box LLM to attempt the data science challenges in KramaBench. DS-Guru has three variants. No-context: The LLM is invoked one-shot with the problem description and the names and paths of the files from the data lake, without any file contents. One-shot: The LLM is invoked one-shot with the problem description and sample snippets from each data file. Few-shot: The LLM is first invoked once with the task description and sample snippets, then re-invoked few-shot with execution results and error messages from the pipeline it implemented in the previous shot. With all variants, DS-Guru instructs the LLM to decompose the task into simpler tasks before attempting to implement each task provide the concrete pipeline implementation along with the answer.
@@ 3 Experimental Setup
We accessed all LLMs in different systems via OpenAI and Together APIs; pipelines generated by systems are executed locally. Whenever possible, we repeated all experiments three times and report average score and standard deviations across the three runs. For some results, marked with an asterisk, we could not perform repeated experiments - due to high API costs or to the backend models being discontinued during the experiments.
DS-Guru: We combine each of the three variants (as in Subsection 2.3) of DS-Guru with six LLMs: GPT-o3, GPT-4o, Claude-3.5-Sonnet, Llama3.3, Deepseek-R1-70B, and Qwen2.5-Coder-32B (OpenAI, 2025b; 2024; Anthropic, 2024; Meta AI, 2025; DeepSeek-AI et al., 2025; Hui et al., 2024), totaling to 18 concrete DS-Guru implementations.
smolagents Deep Research (smolagents DR): We use Hugging Face smolagents (Hugging Face, 2025) to evaluate deep research-style agentic systems on our benchmark under both single-agent and multi-agent settings. On the single agent setting (smolagents-single DR), smolagent’s official single-agent deep research implementation using code actions (69), we report results with two different LLMs as the LLM: GPT-o3 and Claude-3.7-Sonnet. For the multi-agent setting, we implemented two representative architectures also using smolagents: (1) smolagents Reflexion (Shinn et al., 2023) which follows an actor evaluator reflection agent loop. (2) smolagents PDT which follows a Planner and Task Decomposer Tool Executor workflow (PDT), inspired by Fan et al. (2025). We provide more details in Appendix C. We view these systems representative of open-source “deep research” projects (e.g., Alibaba’s Academy (2025)).
Closed-Source Deep Research Systems: OpenAI Deep Research (OpenAI DR, OpenAI (2025a)) and Gemini Pro-2.5 Agentic Mode (Gemini Agentic) (Google DeepMind, 2025; Google, 2025) were evaluated manually through their web interfaces under the end-to-end automation setting. We made best efforts instructing them not to search online. However, this restriction was not enforceable.
@@
| smolagents DR | Agentic retrieval | Full, Trimmed, Oracle | Single-agent ReAct loop | Off |
| smolagents Reflexion | Agentic retrieval | Full, Trimmed, Oracle | Multi-agent | Off |
| smolagents PDT | Agentic retrieval | Full, Trimmed, Oracle | Multi-agent | Off |
@@
Human Baseline. We conducted a small-scale human study in which 9 data-science practitioners solved KramaBench under the same conditions and requirements as LLM systems under test in the end-to-end, full input setting. Details can be found in Appendix F.
We evaluated six different systems, which differ in four important ways.
@@
Control flow describes whether a system has fully structured loops or an agentic workflow where agents decide the future courses of actions. Internet access describes whether systems have web search capabilities.
Cost of evaluation. For DS-Guru (few-shot, GPT-o3), evaluating end-to-end answers, pipeline design, and sub-task implementation took 4,501, 116,805, and 10,358 tokens respectively.
@@ 4 Results and Takeaways
| Archeology | Astronomy | Biomedical | Environment | Legal | Wildfire | Overall | ||
@@
| Human baseline* | 58.33 | 70.00 | 100.00 | 76.90 | 86.67 | 66.20 | 76.75 | |
| DS-Guru no-context | GPT-o3 | 16.67 ± 0.00 | 0.00 ± 0.00 | 0.10 ± 0.14 | 5.00 ± 4.08 | 0.00 ± 0.00 | 14.18 ± 4.28 | 5.87 ± 0.71 |
| GPT-4o | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | |
| Claude-3.7 | 2.78 ± 3.93 | 3.33 ± 4.71 | 0.00 ± 0.00 | 0.00 ± 0.00 | 1.11 ± 1.57 | 4.38 ± 3.92 | 1.88 ± 0.81 | |
| DS-Guru one-shot | GPT-o3 | 19.44 ± 3.93 | 0.00 ± 0.00 | 0.53 ± 0.75 | 16.11 ± 9.26 | 10.00 ± 0.00 | 40.01 ± 4.22 | 16.67 ± 2.93 |
| GPT-4o | 8.33 ± 6.80 | 6.67 ± 4.71 | 3.70 ± 0.00 | 0.00 ± 0.00 | 4.41 ± 2.17 | 13.69 ± 3.06 | 6.08 ± 1.25 | |
| Claude-3.7 | 2.78 ± 3.93 | 0.00 ± 0.00 | 0.00 ± 0.00 | 3.33 ± 2.36 | 0.00 ± 0.00 | 1.59 ± 2.24 | 1.31 ± 1.22 | |
| DS-Guru few-shot | GPT-o3 | 13.89 ± 3.93 | 0.00 ± 0.00 | 0.10 ± 0.14 | 49.56 ± 0.87 | 9.26 ± 2.92 | 52.92 ± 0.72 | 24.98 ± 1.25 |
| GPT-4o | 13.89 ± 3.93 | 3.33 ± 4.71 | 0.00 ± 0.00 | 0.00 ± 0.00 | 4.84 ± 3.16 | 25.69 ± 1.15 | 8.67 ± 1.48 | |
| Claude-3.7 | 5.56 ± 3.93 | 0.00 ± 0.00 | 0.00 ± 0.00 | 15.00 ± 4.08 | 8.89 ± 6.29 | 10.09 ± 9.67 | 8.29 ± 1.12 | |
| smolagents DR | GPT-o3 | 33.33 ± 11.79 | 23.33 ± 9.43 | 37.04 ± 10.48 | 31.33 ± 8.96 | 17.78 ± 8.31 | 35.23 ± 11.46 | 28.07 ± 8.80 |
| Claude-3.7 | 44.44 ± 3.93 | 50.00 ± 8.16 | 38.89 ± 7.86 | 60.56 ± 7.74 | 61.23 ± 12.70 | 60.16 ± 7.43 | 55.83 ± 3.41 | |
| smolagents Reflexion | GPT-o3 | 30.56 ± 10.39 | 46.67 ± 4.71 | 29.63 ± 5.24 | 23.89 ± 14.55 | 26.67 ± 14.40 | 35.04 ± 8.36 | 30.53 ± 10.79 |
| Claude-3.7 | 41.67 ± 0.00 | 45.00 ± 5.00 | 50.00 ± 27.78 | 56.25 ± 8.75 | 58.33 ± 1.67 | 65.38 ± 0.86 | 55.37 ± 3.36 | |
| smolagents PDT | GPT-o3 | 8.33 ± 6.80 | 0.00 ± 0.00 | 0.00 ± 0.00 | 1.42 ± 1.09 | 4.51 ± 0.84 | 22.01 ± 1.39 | 7.57 ± 1.28 |
| Claude-3.7 | 19.44 ± 7.86 | 6.67 ± 4.71 | 5.56 ± 5.56 | 9.84 ± 2.86 | 6.25 ± 4.95 | 22.58 ± 4.05 | 12.01 ± 1.03 | |
@@
| DS-Guru no-context | GPT-o3 | 11.11 ± 3.93 | 10.00 ± 8.16 | 11.11 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 10.17 ± 0.92 | 5.36 ± 1.20 |
| GPT-4o | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | |
| Claude-3.7 | 4.17 ± 4.17 | 10.00 ± 0.00 | 0.00 ± 0.00 | 1.58 ± 1.58 | 1.67 ± 1.67 | 8.96 ± 0.57 | 4.11 ± 0.79 | |
| DS-Guru one-shot | GPT-o3 | 8.33 ± 6.80 | 16.67 ± 4.71 | 7.51 ± 5.31 | 23.89 ± 6.71 | 15.56 ± 1.57 | 42.80 ± 3.48 | 21.35 ± 0.84 |
| GPT-4o | 11.11 ± 7.86 | 13.33 ± 4.71 | 5.19 ± 7.33 | 12.33 ± 3.30 | 6.67 ± 0.00 | 19.85 ± 3.61 | 11.54 ± 1.47 | |
| Claude-3.7 | 0.00 ± 0.00 | 0.00 ± 0.00 | 0.00 ± 0.00 | 6.67 ± 1.67 | 3.33 ± 3.33 | 4.76 ± 0.00 | 3.27 ± 1.31 | |
| DS-Guru few-shot | GPT-o3 | 16.67 ± 0.00 | 26.67 ± 4.71 | 7.41 ± 5.24 | 50.11 ± 4.09 | 61.03 ± 5.78 | 52.02 ± 3.79 | 43.71 ± 1.94 |
| GPT-4o | 13.89 ± 3.93 | 20.00 ± 0.00 | 3.70 ± 5.24 | 18.44 ± 6.00 | 37.78 ± 4.16 | 36.67 ± 3.53 | 26.20 ± 2.34 | |
| Claude-3.7 | 20.83 ± 4.17 | 25.00 ± 15.00 | 0.00 ± 0.00 | 9.17 ± 9.17 | 31.67 ± 1.67 | 18.17 ± 5.20 | 19.75 ± 3.36 | |
| smolagents DR | GPT-o3 | 27.78 ± 7.86 | 26.67 ± 4.71 | 37.04 ± 5.24 | 24.00 ± 0.00 | 17.78 ± 5.67 | 32.89 ± 10.31 | 25.86 ± 4.57 |
| Claude-3.7 | 45.83 ± 12.50 | 65.00 ± 5.00 | 58.33 ± 2.78 | 46.83 ± 1.50 | 65.00 ± 1.67 | 75.08 ± 3.78 | 60.67 ± 1.23 | |
| smolagents Reflexion | GPT-o3 | 30.56 ± 3.93 | 40.00 ± 0.00 | 44.44 ± 0.00 | 31.44 ± 2.20 | 22.22 ± 8.31 | 39.78 ± 6.14 | 32.33 ± 4.44 |
| Claude-3.7 | 37.50 ± 4.17 | 50.00 ± 0.00 | 83.33 ± 5.56 | 62.67 ± 4.33 | 70.00 ± 10.00 | 64.45 ± 1.39 | 62.81 ± 3.10 | |
| smolagents PDT | GPT-o3 | 8.33 ± 6.80 | 0.00 ± 0.00 | 0.00 ± 0.00 | 3.26 ± 1.30 | 8.62 ± 3.88 | 15.86 ± 5.53 | 7.69 ± 2.69 |
| Claude-3.7 | 16.67 ± 0.00 | 10.00 ± 0.00 | 11.11 ± 0.00 | 11.50 ± 3.50 | 9.24 ± 0.35 | 25.50 ± 5.46 | 14.38 ± 2.15 | |
@@
| DS-GURU few-shot | GPT-o3 | 25.00 ± 6.80 | 0.00 ± 0.00 | 3.70 ± 5.24 | 48.33 ± 8.50 | 51.11 ± 3.14 | 48.87 ± 7.90 | 37.84 ± 3.90 |
| smolagents-single DR | Claude-3.7* | 50.00 | 80.00 | 55.56 | 40.83 | 60.00 | 67.24 | 58.12 |
| OpenAI Online Deep Research* | 40.00 | 33.33 | 44.45 | 61.67 | 50.00 | 67.28 | 52.18 | |
| Google Agentic Gemini-2.5-Pro* | 25.00 | 16.67 | 33.33 | 25.00 | 13.33 | 24.87 | 18.48 | |
Agentic control flows drive the largest performance gains on KramaBench. The smolagents DR variant using Claude-3-7 consistently outperforms DS-Guru across all domains (Table 5), reaching an average overall score compared to the best DS-Guru variant (few-shot, GPT-o3; ). The DS-Guru few-shot variant, which enables the LLM to catch implementation errors, improves over the one shot variant, e.g., using GPT-o3, with a overall improvement. Our detailed studies increased few-shot to 20 iterations yet still showed minor improvements (Appendix Table 13). This indicates that despite the heterogeneity of data operations, the core challenges are not isolated data operation implementation issues, but instead are to (1) explore and fix the design choices of the end-to-end pipeline; (2) iteratively understand the data and schema in a large data lake. Smolagent DR’s agentic control flow helps address these challenges. In the Trimmed setting (max 10 files per call), OpenAI’s own online DeepResearch tool reaches overall, partly due to its web search capability. We refer the reader to Appendix F for detailed analysis of the human baseline., as well as for a detailed analysis of cost and runtimes of each models. In terms of cost, smolagents DR (Claude-3-7) averages 6.10 minutes per task—faster than OpenAI DR (10.35) but more than 10× slower than DS-Guru few-shot (0.76).